# Higher equations of motion at level 2 in Liouville CFT

### v1

@@ 2.3. First pole of $\mathcal{P}_{\alpha}$ and the $(1,2)$-HEM
On the other hand, using the intertwining relation and $\widetilde{\mathbf{S}}_{\alpha_{1,2}}\Psi_{\alpha_{1,2}}=0$ [BW23, Theorem 1.2], we have
$\frac{\partial}{\partial\alpha}_{|\alpha=\alpha_{1,2}}\mathcal{P}_{\alpha}(\mathbf{S}^{0}_{\alpha}\widetilde{\mathbf{S}}^{0}_{\alpha}\mathds{1})=\underset{\alpha\to\alpha_{1,2}}{\lim}\,\frac{1}{\alpha-\alpha_{1,2}}\mathbf{S}_{\alpha}\widetilde{\mathbf{S}}_{\alpha}\Psi_{\alpha}$ $=\underset{\alpha\to\alpha_{1,2}}{\lim}\,\mathbf{S}_{\alpha}\widetilde{\mathbf{S}}_{\alpha}\left(\frac{\Psi_{\alpha}-\Psi_{\alpha_{1,2}}}{\alpha-\alpha_{1,2}}\right)$
$=\mathbf{S}_{\alpha_{1,2}}\widetilde{\mathbf{S}}_{\alpha_{1,2}}\Psi_{\alpha_{1,2}}^{\prime}.$
@@ 2.4. Second pole of $\mathcal{P}_{\alpha}$ and the $(2,1)$-HEM
We have in $e^{-\beta c}\mathcal{D}(\mathcal{Q})$
@@
where $\mathcal{R}_{\alpha}$ is analytic at $\alpha_{2,1}$. Moreover,
@@
(Thus, $\mathcal{I}_{2,(1,1),(1,1)}$ is regular at $\alpha_{2,1}$ for $\gamma>\sqrt{2}$.)
@@
By the intertwining relation and Proposition 2.4, we have
$\mathbf{S}_{\alpha_{2,1}}\widetilde{\mathbf{S}}_{\alpha_{2,1}}\Psi_{\alpha_{2,1}}^{\prime}=\frac{\partial}{\partial\alpha}_{\alpha=\alpha_{2,1}}\mathcal{P}_{\alpha}(\mathbf{S}_{\alpha}^{0}\widetilde{\mathbf{S}}_{\alpha}^{0}\mathds{1})$ $=\underset{\alpha\to\alpha_{2,1}}{\lim}\,\frac{1}{\alpha-\alpha_{2,1}}\mathcal{P}_{\alpha}(\mathbf{S}_{\alpha}^{0}\widetilde{\mathbf{S}}_{\alpha}^{0}\mathds{1})$
@@
The remainder of Section 2.4 is devoted to the proof of Proposition 2.4. First, we show that the residue of $\mathcal{I}_{2,(1,1),(1,1)}$ at $\alpha_{2,1}$ factorises into the primary field $\Psi_{\alpha_{-1,2}}$ and a Dotsenko-Fateev integral. This is done thanks to the fusion estimate of Proposition 2.6. The residue of the Dotsenko-Fateev integral can be computed explicitly (Lemma 2.5), and we get (2.4). Finally, Proposition 2.7 establishes the expression of (2.3) for the meromorphic continuation, using a method inspired by [BW23]. Our method for the computation of the scalar multiple of $\Psi_{\alpha_{-2,1}}$ differs (and is somewhat more natural) than the one proposed in [Zam04]: we do purely free field computations (requiring only the exact value of the Dotsenko-Fateev integral), while [Zam04, Section 5] makes a detour through Liouville (using the DOZZ formula and non-trivial properties of the $\Upsilon$-function and Virasoro representation theory) before coming back to the free field.
2.4.1. Residues of Dotsenko-Fateev integrals
@@
$J_{1}(\alpha):=\int_{\mathbb{D}^{2}}|w_{1}|^{-\gamma\alpha}|w_{2}|^{-\gamma\alpha}|w_{1}-w_{2}|^{-\gamma^{2}}\frac{|\mathrm{d}w_{1}|^{2}}{|w_{1}|^{2}}\frac{|\mathrm{d}w_{2}|^{2}}{|w_{2}|^{2}};$
$J_{2}(\alpha):=\int_{\mathbb{D}^{2}}|w_{1}|^{-\gamma\alpha}|w_{2}|^{-\gamma\alpha}|w_{1}-w_{2}|^{-\gamma^{2}}\frac{\bar{w}_{2}-\bar{w}_{1}}{w_{2}-w_{1}}\frac{|\mathrm{d}w_{1}|^{2}}{w_{1}\bar{w}_{1}^{2}}\frac{|\mathrm{d}w_{2}|^{2}}{\bar{w}_{2}};$
$J_{3}(\alpha):=\int_{\mathbb{D}^{2}}|w_{1}|^{-\gamma\alpha}|w_{2}|^{-\gamma\alpha}|w_{1}-w_{2}|^{-\gamma^{2}}\frac{\bar{w}_{2}-\bar{w}_{1}}{w_{2}-w_{1}}\frac{|\mathrm{d}w_{1}|^{2}}{w_{1}\bar{w}_{1}}\frac{|\mathrm{d}w_{2}|^{2}}{\bar{w}_{2}^{2}}.$
The integrals $J_{1}$ and $J_{3}$ are absolutely convergent for $\mathrm{Re}(\alpha)<\alpha_{2,1}$. The integral $J_{2}$ is absolutely convergent for $\mathrm{Re}(\alpha)<-\frac{1}{\gamma}$.
@@
The functions $J_{2}$, $J_{3}$ are regular at $\alpha_{2,1}$.
@@
which is well-defined and analytic for $\mathrm{Re}(\alpha)<\alpha_{2,1}$ and suitables values of $\beta$ as follows. If $D=\mathbb{D}$, then $J_{D}^{\beta}$ is well defined and analytic for $\beta$ in an open neighbourhood of 0, and the quantity we want to compute is $\underset{\alpha=\alpha_{2,1}}{\mathrm{Res}}\,J_{\mathbb{D}}^{\beta=0}(\alpha)$. If $D=\mathbb{C}$, we need $\mathrm{Re}(\beta)>-\mathrm{Re}(\alpha+\frac{\gamma}{2})$ for the integrability at $\infty$. In the region of absolute convergence, $J_{\mathbb{C}}^{\beta}(\alpha)$ is the Dotsenko-Fateev integral of Appendix A, whose meromorphic continuation is given by (A.3):
@@
The last integral is regular at $\alpha=\alpha_{2,1}$ for all $\beta$ in a neighbourhood of 0 (indeed, the first pole is at $\alpha=0$ uniformly in $\beta$). Moreover, the last formula defines the meromorphic extension of $J_{\mathbb{D}}^{\beta}$ in a neighbourhood of $\alpha_{2,1}$. Let $A_{\beta}:=\underset{\alpha=\alpha_{2,1}}{\mathrm{Res}}\,J_{\mathbb{D}}^{\beta}(\alpha)$; the quantity we want to compute is $A_{0}$. Writing the Laurent expansions around $\alpha_{2,1}$ gives
$J_{\mathbb{C}}^{\beta}(\alpha)$ $=\frac{A_{\beta}}{\alpha-\alpha_{2,1}}-\frac{A_{\beta}}{\alpha-\alpha_{2,1}+\beta}+f^{\beta}(\alpha)=\frac{\beta A_{\beta}}{(\alpha-\alpha_{2,1})(\alpha-\alpha_{2,1}+\beta)}(1+o(1)),$
where $f^{\beta}(\alpha)$ is analytic for $(\alpha,\beta)$ in a neighbourhood of $(\alpha_{2,1},0)$. Comparing with (2.5), we find
$A_{0}=-\pi\frac{2}{\gamma}\left(\frac{\Gamma(\frac{\gamma^{2}}{4})\Gamma(1-\frac{\gamma^{2}}{2})}{\Gamma(1-\frac{\gamma^{2}}{4})}\right)^{2}\sin(\pi\frac{\gamma^{2}}{2})=-\pi^{2}\frac{2}{\gamma}\left(\frac{\Gamma(\frac{\gamma^{2}}{4})}{\Gamma(1-\frac{\gamma^{2}}{4})}\right)^{2}\frac{\Gamma(1-\frac{\gamma^{2}}{2})}{\Gamma(\frac{\gamma^{2}}{2})}.$
Regularity of $J_{2}$.
For $\mathrm{Re}(\alpha)<\alpha_{2,1}$, we have by integration by parts
$\int_{\mathbb{D}^{2}}|w_{1}|^{-\gamma\alpha}|w_{2}|^{-\gamma\alpha}|w_{1}-w_{2}|^{-\gamma^{2}}\frac{|\mathrm{d}w_{1}|^{2}}{w_{1}\bar{w}_{1}^{2}}\frac{|\mathrm{d}w_{2}|^{2}}{\bar{w}_{2}}$ $=\frac{\frac{\gamma\alpha}{2}(\frac{\gamma\alpha}{2}+1)}{(1-\frac{\gamma^{2}}{2})^{2}}\int_{\mathbb{D}^{2}}|w_{1}|^{-\gamma\alpha}|w_{2}|^{-\gamma\alpha}|w_{1}-w_{2}|^{2-\gamma^{2}}\frac{|\mathrm{d}w_{1}|^{2}}{w_{1}\bar{w}_{1}^{2}}\frac{|\mathrm{d}w_{2}|^{2}}{w_{2}\bar{w}_{2}^{2}}$
$-\frac{1}{2i}\frac{\frac{\gamma\alpha}{2}}{(1-\frac{\gamma^{2}}{2})^{2}}\int_{\mathbb{D}}\int_{\mathbb{S}^{1}}|w_{1}|^{-\gamma\alpha}|w_{1}-w_{2}|^{2-\gamma^{2}}\mathrm{d}w_{2}\frac{|\mathrm{d}w_{1}|^{2}}{w_{1}\bar{w}_{1}^{2}}$
$+\frac{1}{2i}\frac{1}{1-\frac{\gamma^{2}}{2}}\int_{\mathbb{D}}\int_{\mathbb{S}^{1}}|w_{1}|^{-\gamma\alpha}\frac{|w_{1}-w_{2}|^{2-\gamma^{2}}}{\bar{w}_{2}-\bar{w}_{1}}\frac{\mathrm{d}\bar{w}_{2}}{\bar{w}_{2}}\frac{|\mathrm{d}w_{1}|^{2}}{w_{1}\bar{w}_{1}^{2}}.$
The two boundary integrals are absolutely convergent and analytic in the neighbourhood of $\alpha_{2,1}$. It remains to treat the bulk integral, which we denote $\widetilde{J}_{2}(\alpha)$. Similarly to the previous paragraph, we consider the auxiliary function, for $D\in\{\mathbb{D},\mathbb{C}\}$:
$\widetilde{J}_{D,2}^{\beta}(\alpha):=\int_{D^{2}}|w_{1}|^{-\gamma\alpha-2}|w_{2}|^{-\gamma\alpha-2}|1-w_{1}|^{-\gamma\beta}|1-w_{2}|^{-\gamma\beta}|w_{1}-w_{2}|^{2-\gamma^{2}}\frac{|\mathrm{d}w_{1}|^{2}}{\bar{w}_{1}}\frac{|\mathrm{d}w_{2}|^{2}}{\bar{w}_{2}}.$
From Appendix A, we have
$J_{\mathbb{C},2}^{\beta}(\alpha)$ $=S_{2,2}(-\frac{\gamma\alpha}{2},1-\frac{\gamma\beta}{2},-\frac{\gamma^{2}}{4})S_{2,2}(-\frac{\gamma\alpha}{2}-1,1-\frac{\gamma\beta}{2},1-\frac{\gamma^{2}}{4})$
×$\quad\times\frac{\sin(\pi\frac{\gamma\alpha}{2})\sin(\pi\frac{\beta\gamma}{2})\sin\pi\frac{\gamma}{2}(\alpha+\gamma)\sin\pi\frac{\gamma}{2}(\beta+\gamma)\sin(\pi\frac{\gamma^{2}}{2})}{\sin\pi\frac{\gamma}{2}(\alpha+\beta+\frac{\gamma}{2})\sin\pi\frac{\gamma}{2}(\alpha+\beta+\gamma)\sin(\pi\frac{\gamma^{2}}{4})}.$
From this expression, we can apply the same method as for $J_{1}$, and we find that $\widetilde{J}_{\mathbb{D},2}^{0}$ has a simple pole at $\alpha=\alpha_{2,1}$. Since $J_{2}$ is proportional to $(\alpha-\alpha_{2,1})\widetilde{J}_{2}$ (up to analytic terms around $\alpha_{2,1}$), we deduce that $J_{2}$ is analytic at $\alpha_{2,1}$.
Regularity of $J_{3}$.
This can be handled similarly to $J_{2}$, and we omit the proof. ∎

### v2

@@ 2.3. First pole of $\mathcal{P}_{\alpha}$ and the $(1,2)$-HEM
On the other hand, [BGK+24a, Theorem 4.5] states that the Liouville descendant states are obtained from the primary state $\Psi_{\alpha}$ by acting with the Virasoro generators of the Liouville theory: in other words, the Poisson operator intertwines the Virasoro algebras of the free field theory and the Liouville theory in the sense that $\mathcal{P}_{\alpha}\circ\mathbf{L}_{n}^{0,\alpha}=\mathbf{L}_{n}^{\alpha}\circ\mathcal{P}_{\alpha}$, with $\mathbf{L}_{n}^{\alpha}$ the operator $\mathbf{L}_{n}$ acting in the module $\mathcal{W}_{\alpha}$ defined in [BGK+24a, Section 4.4]. The same also holds with the representation $(\widetilde{\mathbf{L}}_{n})_{n\in\mathbb{Z}}$. Combining with $\widetilde{\mathbf{S}}_{\alpha_{1,2}}\Psi_{\alpha_{1,2}}=0$ [BW23, Theorem 1.2], we have
$\frac{\partial}{\partial\alpha}_{|\alpha=\alpha_{1,2}}\mathcal{P}_{\alpha}(\mathbf{S}^{0}_{\alpha}\widetilde{\mathbf{S}}^{0}_{\alpha}\mathds{1})=\underset{\alpha\to\alpha_{1,2}}{\lim}\,\frac{1}{\alpha-\alpha_{1,2}}\mathbf{S}_{\alpha}\widetilde{\mathbf{S}}_{\alpha}\Psi_{\alpha}=\underset{\alpha\to\alpha_{1,2}}{\lim}\,\mathbf{S}_{\alpha}\widetilde{\mathbf{S}}_{\alpha}\left(\frac{\Psi_{\alpha}-\Psi_{\alpha_{1,2}}}{\alpha-\alpha_{1,2}}\right)=\mathbf{S}_{\alpha_{1,2}}\widetilde{\mathbf{S}}_{\alpha_{1,2}}\Psi_{\alpha_{1,2}}^{\prime}.$
@@ 2.4. Second pole of $\mathcal{P}_{\alpha}$ and the $(2,1)$-HEM
We have in $e^{-\beta c}\mathcal{D}(\mathcal{Q})$ for $\alpha\in(-\infty,\alpha_{1,2})$
@@
where $\mathcal{R}_{\alpha}$ is analytic up to a neighbourhood of $\alpha_{2,1}$. Moreover, $\mathcal{I}_{2,(1,1),(1,1)}$ is meromorphic in a neighbourhood of $(-\infty,\alpha_{2,1})$, with
@@
In particular, $\mathcal{I}_{2,(1,1),(1,1)}$ is regular at $\alpha_{2,1}$ if $\gamma>\sqrt{2}$.
@@
As in the previous proof, the intertwining relation of [BGK+24a, Theorem 4.5] and Proposition 2.4 give
$\mathbf{S}_{\alpha_{2,1}}\widetilde{\mathbf{S}}_{\alpha_{2,1}}\Psi_{\alpha_{2,1}}^{\prime}=\frac{\partial}{\partial\alpha}_{|\alpha=\alpha_{2,1}}\mathcal{P}_{\alpha}(\mathbf{S}_{\alpha}^{0}\widetilde{\mathbf{S}}_{\alpha}^{0}\mathds{1})$ $=\underset{\alpha\to\alpha_{2,1}}{\lim}\,\frac{1}{\alpha-\alpha_{2,1}}\mathcal{P}_{\alpha}(\mathbf{S}_{\alpha}^{0}\widetilde{\mathbf{S}}_{\alpha}^{0}\mathds{1})$
@@
The remainder of Section 2.4 is devoted to the proof of Proposition 2.4. First, we show that the residue of $\mathcal{I}_{2,(1,1),(1,1)}$ at $\alpha_{2,1}$ factorises into the primary field $\Psi_{\alpha_{-1,2}}$ and a Dotsenko–Fateev integral. This is done thanks to the fusion estimate of Proposition 2.6. The residue of the Dotsenko–Fateev integral can be computed explicitly (Lemma 2.5), and we get (2.8). Finally, Proposition 2.7 establishes the expression of (2.7) for the meromorphic continuation, using a method inspired by [BW23]. Our method for the computation of the scalar multiple of $\Psi_{\alpha_{-2,1}}$ differs (and is somewhat more natural) from the one proposed in [ZAM04]: we do purely free field computations (requiring only the exact value of the Dotsenko–Fateev integral), while [ZAM04, Section 5] makes a detour through Liouville (using the DOZZ formula and non-trivial properties of the $\Upsilon$-function and Virasoro representation theory) before coming back to the free field.
2.4.1. Residues of Dotsenko–Fateev integrals
@@
$J_{1}(\alpha)=\int_{\mathbb{D}^{2}}|w_{1}|^{-\gamma\alpha}|w_{2}|^{-\gamma\alpha}|w_{1}-w_{2}|^{-\gamma^{2}}\frac{|\mathrm{d}w_{1}|^{2}}{|w_{1}|^{2}}\frac{|\mathrm{d}w_{2}|^{2}}{|w_{2}|^{2}};$ (2.9)
$J_{2}(\alpha)=\int_{\mathbb{D}^{2}}|w_{1}|^{-\gamma\alpha}|w_{2}|^{-\gamma\alpha}|w_{1}-w_{2}|^{-\gamma^{2}}\frac{\bar{w}_{2}-\bar{w}_{1}}{w_{2}-w_{1}}\frac{|\mathrm{d}w_{1}|^{2}}{w_{1}\bar{w}_{1}^{2}}\frac{|\mathrm{d}w_{2}|^{2}}{\bar{w}_{2}};$
$J_{3}(\alpha)=\int_{\mathbb{D}^{2}}|w_{1}|^{-\gamma\alpha}|w_{2}|^{-\gamma\alpha}|w_{1}-w_{2}|^{-\gamma^{2}}\frac{\bar{w}_{2}-\bar{w}_{1}}{w_{2}-w_{1}}\frac{|\mathrm{d}w_{1}|^{2}}{w_{1}\bar{w}_{1}}\frac{|\mathrm{d}w_{2}|^{2}}{\bar{w}_{2}^{2}}.$
The integrals $J_{1}$ and $J_{3}$ are absolutely convergent for $\mathrm{Re}(\alpha)<\alpha_{2,1}=-\frac{\gamma}{2}$. The integral $J_{2}$ is absolutely convergent for $\mathrm{Re}(\alpha)<-\frac{1}{\gamma}$.
@@
The function $J_{2}+J_{3}$ is regular at $\alpha_{2,1}$.
@@
which is well-defined and analytic for $\mathrm{Re}(\alpha)\in(\alpha_{2,1},0)$ and $\mathrm{Re}(\beta)\in(\alpha_{2,1}-\alpha,\alpha_{2,1}+\frac{2}{\gamma})$ if $D=\mathbb{C}$. If $D=\mathbb{D}$, we have the same bound on $\alpha$, but $J_{\mathbb{D}}$ is actually analytic all the way to $\beta$ in a neighbourhood of 0 (in fact, we want to compute $\mathrm{Res}_{\alpha=\alpha_{2,1}}J_{\mathbb{D}}^{\beta=0}(\alpha)$). In the region of absolute convergence, $J_{\mathbb{C}}^{\beta}(\alpha)$ is the Dotsenko–Fateev integral of Appendix A, whose meromorphic continuation is given by (A.3):
@@
The last integral is regular at $\alpha=\alpha_{2,1}$ for all $\beta$ in a neighbourhood of 0 (indeed, the first pole is at $\alpha=0$ uniformly in $\beta$, and $\gamma<\sqrt{2}$). Moreover, since $J_{\mathbb{D}}^{\beta}$ has a pole at $\alpha_{2,1}$, the second term of the previous displayed equation has a pole at $\alpha_{2,1}-\beta$. Now, note that $J_{\mathbb{D}}^{\beta}$ converges uniformly on compacts away from $\alpha_{2,1}$ as $\beta\to 0$, so $J_{\mathbb{D}}^{\beta}$ does not have a pole at $\alpha_{2,1}-\beta$ for small enough $\beta>0$. It follows that $\mathrm{Res}_{\alpha=\alpha_{2,1}}J_{\mathbb{D}}^{\beta}(\alpha)=\mathrm{Res}_{\alpha=\alpha_{2,1}}J_{\mathbb{C}}^{\beta}(\alpha)$ for sufficiently small $\mathrm{Re}(\beta)>0$. Taking $\beta\to 0$ gives
$\mathrm{Res}_{\alpha=\alpha_{2,1}}\,J_{\mathbb{D}}(\alpha)=-\frac{2\pi}{\gamma}\left(\frac{\Gamma(\frac{\gamma^{2}}{4})\Gamma(1-\frac{\gamma^{2}}{2})}{\Gamma(1-\frac{\gamma^{2}}{4})}\right)^{2}\sin(\pi\frac{\gamma^{2}}{2})=-\pi^{2}\frac{2}{\gamma}\left(\frac{\Gamma(\frac{\gamma^{2}}{4})}{\Gamma(1-\frac{\gamma^{2}}{4})}\right)^{2}\frac{\Gamma(1-\frac{\gamma^{2}}{2})}{\Gamma(\frac{\gamma^{2}}{2})}.$
Regularity of $J_{2}+J_{3}$.
Fix $\alpha$ with $\mathrm{Re}(\alpha)<\alpha_{2,1}$.
For $J_{2}$, we do an integration by parts with respect to $\partial_{\bar{w}_{2}}((w_{1}-w_{2})^{-\frac{\gamma^{2}}{2}-1}(\bar{w}_{1}-\bar{w}_{2})^{-\frac{\gamma^{2}}{2}+2})$
$J_{2}=$ $-\frac{-\frac{\gamma\alpha}{2}-1}{\frac{\gamma^{2}}{2}-2}\int_{\mathbb{D}^{2}}|w_{1}|^{-\gamma\alpha}|w_{2}|^{-\gamma\alpha}|w_{1}-w_{2}|^{-\gamma^{2}}\frac{(\bar{w}_{1}-\bar{w}_{2})^{2}}{w_{1}-w_{2}}\frac{|\mathrm{d}w_{1}|^{2}}{w_{1}\bar{w}_{1}^{2}}\frac{|\mathrm{d}w_{2}|^{2}}{\bar{w}_{2}^{2}}$
$\quad+\frac{i}{2}\frac{1}{\frac{\gamma^{2}}{2}-2}\int_{\mathbb{S}^{1}}\int_{\mathbb{D}}|w_{2}|^{-\gamma\alpha}|w_{1}-w_{2}|^{-\gamma^{2}}\frac{(\bar{w}_{1}-\bar{w}_{2})^{2}}{w_{1}-w_{2}}\frac{|\mathrm{d}w_{1}|^{2}}{\bar{w}_{1}}\frac{\mathrm{d}w_{2}}{\bar{w}_{2}}.$
For $J_{3}$, we do integration by parts with respect to $\partial_{\bar{w}_{1}}((w_{1}-w_{2})^{-\frac{\gamma^{2}}{2}-1}(\bar{w}_{1}-\bar{w}_{2})^{-\frac{\gamma^{2}}{2}+2})$
$J_{3}$ $=-\frac{-\frac{\gamma\alpha}{2}-1}{2-\frac{\gamma^{2}}{2}}\int_{\mathbb{D}^{2}}|w_{1}|^{-\gamma\alpha}|w_{2}|^{-\gamma\alpha}|w_{1}-w_{2}|^{-\gamma^{2}}\frac{(\bar{w}_{1}-\bar{w}_{2})^{2}}{w_{1}-w_{2}}\frac{|\mathrm{d}w_{1}|^{2}}{w_{1}\bar{w}_{1}^{2}}\frac{|\mathrm{d}w_{2}|^{2}}{\bar{w}_{2}^{2}}$
$\quad+\frac{i}{2}\frac{1}{2-\frac{\gamma^{2}}{2}}\int_{\mathbb{D}}\int_{\mathbb{S}^{1}}|w_{2}|^{-\gamma\alpha}|w_{1}-w_{2}|^{-\gamma^{2}}\frac{(\bar{w}_{1}-\bar{w}_{2})^{2}}{w_{1}-w_{2}}\mathrm{d}w_{1}\frac{|\mathrm{d}w_{2}|^{2}}{\bar{w}_{2}^{2}}.$
Then we see the bulk terms of $J_{2}$ and $J_{3}$ cancel, and the boundary terms are absolutely convergent at $\alpha_{2,1}$. ∎